# Cascade Speculative Drafting for Even Faster LLM Inference

### v4

@@ 5.3 Experimental Results
Comparison to Baselines Table 2 presents the experimental results. In two settings of SWI, Cascade Speculative Decoding has outperformed the speculative decoding algorithm. For GSM8K, Cascade Speculative Decoding achieved a maximum additional speedup of 44% over the fastest speculative algorithm; for MMLU, the maximum additional speedup improvement over speculative decoding is 81%. Overall, CS Decoding achieves a 3-4x speedup over autoregressive generation, establishing itself as a faster algorithm over the baselines.
Effectiveness of MaG When comparing CS Drafting with one neural model and MaG against the fastest speculative decoding setup, we found that CS Drafting with one neural model gained up to a 70% speedup on MMLU and a 32% speedup on GSM8K. Notably, the MaG algorithm only involves a bigram model with as many parameters as the size of the tokenizer, making its memory cost negligible. Notably, the speedup gained using CS Drafting with one neural model involves no additional deployment overhead while reducing both latency and computational cost, making it a superior choice over speculative decoding.
@@
Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup $k_{11}$
GSM8K Auto. - 1 -
GSM8K S Decoding 160m $2.48$ 12
GSM8K CS Drafting 160m, mag 2.86 15
MMLU Auto. - 1 -
MMLU S Decoding 160m $2.12$ 10
MMLU CS Drafting 160m, mag 2.64 13
Table 3: The experimental results on LLAMA.
Results of Decoder-only Models As shown in Table 3, the results for decoder-only models exhibit similar traits to those observed in the experiments on encoder-decoder models. Therefore, for speculative decoding-based algorithms, we believe there is no distinct difference between an encoder-decoder model and a decoder-only model, aligning with the results in Leviathan et al. (2023).
@@ 6.1 Efficienct Methods for Language Model Inference
In the era of large language models, efficiency during inference becomes a key to model service. To reduce the model inference cost and speed up, several efficient methods have been proposed, including pruning, knowledge distillation and quantization (Treviso et al., 2023). Model pruning takes structured (Xia et al., 2022; Voita et al., 2019) or unstructured (Guo et al., 2021; Chen et al., 2020) methods to remove the redundant model parameters to reduce the storage memory and increase inference speed. Knowledge distillation takes the approach of transferring knowledge from a superior teacher model to a smaller student model (Hinton et al., 2015; Gou et al., 2021). Quantization maps high-precision data representations (e.g. 32 bits) into low-precision ones (e.g. 8 bits) to reduce memory consumption (Bhandare et al., 2019; Schaefer et al., 2023).
@@ 6.2 Speculative Decoding
With the success of Speculative Decoding (Chen et al., 2023; Leviathan et al., 2023) in reducing the large language model inference latency, some recent works have attempted to improve Speculative Decoding by reducing the rejection rate. Zhou et al. (2023) propose using generalized knowledge distillation and achieve a lower rejection rate compared to other knowledge distillation methods. Avoiding an additional draft model, self-drafting is an approach to speculative decoding by reusing part of the target model together with added weight to perform drafting (Zhang et al., 2023; Hooper et al., 2024). Tree attention involves generating multiple candidates during drafting to increase the chance of acceptance (Spector & Re, 2023; Miao et al., 2024). Besides reducing the rejection rate, improving drafting efficiency can also reduce latency. Spector & Re (2023) propose using speculative decoding for drafting, showing similarities to the vertical cascade; however, their method only has two layers of speculative decoding and does not observe the recursive nature of the vertical cascade nor the lenience among draft models, two crucial aspects for the performance of vertical cascade.
@@ 7 Conclusion
In this work, we propose a novel algorithm, CS Drafting, which involves two cascades: the vertical cascade and the horizontal cascade. The vertical cascade eliminates the necessity of autoregressive generation from a neural language model, while the horizontal cascade effectively allocates the cost of drafting tokens at different positions. Our experiments show that CS Drafting achieves up to an 81 percent additional speedup over speculative decoding, while maintaining the same output distribution as the target model. Our work demonstrates that LLM inference can be further sped up by cascades without sacrificing generation quality or requiring additional task-specific training.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Max-Gram Implementation
Appendix B Additonal Results
Figure 3: The probability of acceptance of draft tokens in relation to their positions in a single step of speculative decoding, evaluated on FLAN-T5-small, base, and large models on MMLU.
@@ Appendix C Proof
To achieve the operator, we begin with a single polynomial term of $x^{j}$. Fortunately, given the end result of $\frac{1-\alpha^{j+1}}{(1-\alpha)}$ (Leviathan et al., 2023), the operator $T_{\alpha}(f(x))=\frac{1-\alpha f(\alpha)}{(1-\alpha)}$ will convert $x^{j}$ to $\frac{1-\alpha^{j+1}}{(1-\alpha)}$. In addition, due to the linearity of the operator, this can be extended to any polynomial. Therefore, we achieved the desired operator to map a probability-generating function into the desired expectation.

### v5

@@ 5.2 Experimental Results
Table 2 presents the main experimental results. In two settings of SWI, Cascade Speculative Drafting has outperformed the speculative decoding algorithm. For GSM8K, CS Drafting achieved a maximum additional speedup of 44% over the fastest speculative algorithm; for MMLU, the maximum additional speedup improvement over speculative decoding is 81%.
Effectiveness of MaG When comparing CS Drafting with one neural model and MaG against the fastest speculative decoding setup, we found that CS Drafting with one neural model gained up to a 70% speedup on MMLU and a 32% speedup on GSM8K. Notably, the MaG algorithm only involves a bigram model with parameters equal to the tokenizer size, making its memory cost negligible." In addition, the speedup gained using CS Drafting with one neural model involves no additional deployment overhead while reducing both latency and computational cost, making it a superior choice over speculative decoding.
@@
Algorithm GSM8K Walltime (Tokens/s) MMLU Walltime (Tokens/s)
Autoregressive 33.34 33.11
S Decoding 44.31 43.65
CS Drafting 56.72 55.60
Medusa 61.87 56.19
CS Drafting + Tree Attention 63.81 63.37
Table 3: The experimental results on Vicuna-7B.
Results of Decoder-only Models As shown in Table 3, CS Drafting achieves a significant walltime improvement over speculative decoding. Moreover, CS Drafting exceeds Medusa when combined with tree attention [15, 3]. This suggests that CS Drafting can be integrated with other efficient designs for speculative decoding to further accelerate inference. We leave the exploration of more advanced combinations as future work.
Ablation Study We perform a hyperparameter study on $K_{00}$, the hyperparameter with the greatest effect on our experiments. As shown in Table 4, the performance of CS Drafting only decreases slightly when this hyperparameter is sub-optimal. Therefore, end users who do not require maximum performance can use a simple setup to achieve near-optimal performance with CS Drafting. Furthermore, we conduct an ablation study by removing the horizontal cascade. On GSM8K, with Vicuna-7B and $K_{00}=1$, this results in a performance drop from 56.16 to 53.55.
$K_{00}$ Walltime (tokens/s)
1 56.16
2 55.51
3 53.73
Table 4: Results on GSM8K with Vicuna-7B under different generation length limits.
@@ 6.1 Efficienct Methods for Language Model Inference
In the era of large language models, efficiency during inference becomes a key to model service. To reduce the model inference cost and speed up, several efficient methods have been proposed, including pruning, knowledge distillation and quantization [21]. Model pruning takes structured [26, 22] or unstructured [9, 5] methods to remove the redundant model parameters to reduce the storage memory and increase inference speed. Knowledge distillation takes the approach of transferring knowledge from a superior teacher model to a smaller student model [11, 8]. Quantization maps high-precision data representations (e.g. 32 bits) into low-precision ones (e.g. 8 bits) to reduce memory consumption [1, 18].
@@ 6.2 Speculative Decoding
With the success of Speculative Decoding [4, 14] in reducing the large language model inference latency, some recent works have attempted to improve Speculative Decoding by reducing the rejection rate. Zhou et al. [29] propose using generalized knowledge distillation and achieve a lower rejection rate compared to other knowledge distillation methods. Avoiding an additional draft model, self-drafting is an approach to speculative decoding by reusing part of the target model together with added weight to perform drafting [27, 12]. Tree attention involves generating multiple candidates during drafting to increase the chance of acceptance [19, 15]. Besides reducing the rejection rate, improving drafting efficiency can also reduce latency. Spector et al. [19] propose using speculative decoding for drafting, showing similarities to the vertical cascade; however, their method only has two layers of speculative decoding and does not observe the recursive nature of the vertical cascade nor the lenience among draft models, two crucial aspects for the performance of vertical cascade.
@@ 7 Conclusion
In this work, we propose a novel algorithm, CS Drafting, which involves two cascades: the vertical cascade and the horizontal cascade. The vertical cascade eliminates the necessity of autoregressive generation from a neural language model, while the horizontal cascade effectively allocates the cost of drafting tokens at different positions. CS Drafting achieves additional speedup over baselines in various settings while maintaining the same output distribution as the target model.
8 Limitations
Our experiments demonstrate strong performance for our methods. However, it is possible, though unlikely, that the outcome might differ on a system with different hardware configurations. To account for this, we report both standardized walltime improvement and raw walltime for a more robust evaluation. Additionally, we provide theoretical analyses to justify the improvements, which is independent of the hardware configurations.
Acknowledgement
This material is based upon work supported by the National Science Foundation IIS 16-19302 and IIS 16-33755, Zhejiang University ZJU Research 083650, IBM-Illinois Center for Cognitive Computing Systems Research (C3SR) and IBM-Illinois Discovery Accelerator Institute (IIDAI), grants from eBay and Microsoft Azure, UIUC OVCR CCIL Planning Grant 434S34, UIUC CSBS Small Grant 434C8U, and UIUC New Frontiers Initiative. Any opinions, findings, conclusions, or recommendations expressed in this publication are those of the author(s) and do not necessarily reflect the views of the funding agencies.
@@ Appendix A Max-Gram Implementation
Appendix B Hyperparameters
To reduce the number of hyperparameters to tune, we use MaG to generate 10 tokens at once, as it is rare for more than 10 tokens to be accepted with the exception when CS Drafting is combined with tree attention. We do not use lenience when the reviewer is $\mathcal{M}_{t}$ to ensure the output distribution does not change. We also avoid lenience between MaG and its reviewer, since there is still a significant performance gap between MaG and a neural model. With these constraints, we are left with at most four hyperparameters: $k_{11}$, $k_{12}$, $k_{22}$, and $l$. For the CS Drafting step where the target model is the reviewer, $k_{11}$ and $k_{12}$ are used. $k_{21}$ and $l$ are used in the step where $\mathcal{M}_{d_{1}}$ is the reviewer. The results of experiments on encoder-decoder models with hyperparameters are shown in Table 5.
When performing experiments with the decoder-only model, we fixed the hyperparameters of CS Drafting for different datasets to better align with most users who do not perform hyperparameter tuning. The k-matrix for CS Drafting is $[[2,10],[0,10]]$. When adding tree attention, we limit it to only the lead node with the highest probability of having children; the k-matrix is $[[1,3],[0,1]]$ with the number of children for each leading node being 8, while the other nodes have no children.
Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup (MS) $k_{11}$ $k_{12}$ $k_{22}$ $l$
GSM8K Autoregressive - 1 - - - -
[…]
MMLU Autoregressive - 1 - - - -
[…]
Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup (PW) $k_{11}$ $k_{12}$ $k_{22}$ $l$
GSM8K Autoregressive - 1 - - - -
[…]
MMLU Autoregressive - 1 - - - -
[…]
Table 5: The experimental results on FLAN-T5 with hyperparameter details. Speedup (MS) is the standardized walltime improvement with the assumption that the latency of each run of a model is its number of parameters (model size). Speedup (PW) is the SWI with the assumption that the latency of each run of a model is the time cost data reported from previous work [14]. $k_{11}$, $k_{12}$, $k_{22}$, $l$ are the hyperparameters. $k_{11}$ and $k_{12}$ represent the step limitation target model and the draft models, $k_{22}$ is the step limitations between the first and second draft model, and $l$ is lenience as shown in algorithm 1. For speculative decoding, the $k_{11}$ is simply the $k$.
@@ Appendix C Proof
To achieve the operator, we begin with a single polynomial term of $x^{j}$. Fortunately, given the end result of $\frac{1-\alpha^{j+1}}{(1-\alpha)}$ [14], the operator $T_{\alpha}(f(x))=\frac{1-\alpha f(\alpha)}{(1-\alpha)}$ will convert $x^{j}$ to $\frac{1-\alpha^{j+1}}{(1-\alpha)}$. In addition, due to the linearity of the operator, this can be extended to any polynomial. Therefore, we achieved the desired operator to map a probability-generating function into the desired expectation.